OpenStudy (anonymous):

Evaluate dy/(dx) for the equation 2xy=12 at the given point (-6,-1). Round answer to two decimal places.

OpenStudy (amistre64):

2'xy + 2x'y + 2xy' = 0

OpenStudy (amistre64):

0xy + 2y + 2x y' = 0 y + x y' = 0 y' = y/x

OpenStudy (anonymous):

2xy = 12 xy = 6 differentiating both sides with respect to x x' y + x y' = 0 y + x dy/dx = 0 Because derivative of x is 1 and y' = dy/dx put in the value of x and y and get your answer!

OpenStudy (anonymous):

Thanks amistre! :)

OpenStudy (amistre64):

yep, i dropped a negative on mine some :)

OpenStudy (amistre64):

brain says one thing and fingers type another ... :)

OpenStudy (anonymous):

Haha true :)